OpenStudy (anonymous):

41five divided by 3five in the given base

OpenStudy (anonymous):

also have this one? Multiply 322base seven X 5 base seven in the given base

OpenStudy (anonymous):

1111 base two + 1000 base two + 1100 base two=1001 base two

OpenStudy (anonymous):

362baseseven+526base seven in the given base

OpenStudy (ranga):

I will answer the last but one question. I am guessing you may have a typo where the = sign should have been a + How to convert base 2 to decimal? Convert abcd base 2 to decimal abcd base 2 = d + c(2^1) + b(2^2) + a(2^3) Therefore, 1111 base 2 = 1 + 1(2^1) + 1(2^2) + 1(2^3) = 1 + 2 + 4 + 8 = 15 1000 base 2 = 0 + 0(2^1) + 0(2^2) + 1(2^3) = 8 1100 base 2 = 0 + 0(2^1) + 1(2^2) + 1(2^3) = 4 + 8 = 12 1001 base 2 = 1 + 0(2^1) + 0(2^2) + 1(2^3) = 1 + 8 = 9 Add them all up: 15 + 8 + 12 + 9 = 44 (which is a decimal). Convert 44 back to base 2 (also called binary). Keep on dividing 44 by 2 22 R 0 11 R 0 5 R 1 2 R 1 1 R 0 0 R 1 44 base 10 = 101100 base 2